# A Foundation of the Boltzmann Distribution, or Tiltings as the Unique Support-Preserving Convolution Endomorphisms

### v2

@@ Abstract.
The family of Boltzmann distributions is used in statistical mechanics to describe the distribution of states in systems with a given temperature. We give a novel characterization of this family as the unique one satisfying independence for uncoupled systems. The theorem boils down to a statement about endomorphisms of the convolution semigroup of finitely supported probability measures on the natural numbers, or, alternatively, about endomorphisms of the multiplicative semigroup of polynomials with non-negative coefficients.
@@ 1. Introduction
In statistical mechanics, a system can be found in various possible states. A state refers to a complete microscopic description of the system which encodes all physically relevant variables, and in particular determines the system’s energy. [^1] Statistical mechanics serves only as a motivation for the mathematical question we study. For a brief introduction to the relevant physics concepts, see Section 2.1 of [mezard2009information]. An uninterested reader may proceed directly to the “Definitions and Results” section. In an equilibrium with a given temperature, the distribution of states is given by the Boltzmann distribution: The probability that the system has energy $E$ is proportional to the number (or mass) of states with energy $E$ times $\mathrm{e}^{-\beta E}$, where $1/\beta$ is proportional to the temperature of the environment. Formally, if $\mu$ is a probability measure on the reals describing the mass of states at each energy—i.e., for measurable $A\subseteq\mathbb{R}$, the mass of states with energy in $A$ is $\mu(A)$—then the distribution of energy is given by the measure $\Phi_{\beta}[\mu]$ where
@@
In terms of the physics, the first property corresponds to the fact that we cannot find the system at an energy that is not attained at any state, and that any state can be attained. The second property involves products of independent systems. Suppose that $\mu_{1}$ and $\mu_{2}$ describe the distribution of states in two systems. Form a new system whose set of states is the product of the two sets of states, and whose energy in each state is the sum of the two corresponding energies; this corresponds to no interaction between the two subsystems. Then the convolution $\mu_{1}*\mu_{2}$ describes the new system, and (1.1) follows from the assumption that the joint distribution of states is the product measure of the distributions in the two subsystems, or, differently put, the assumption that there is no correlation between systems that do not interact.
@@
Given a subset $S\subseteq\mathbb{R}$ closed with respect to addition, denote by $\mathrm{Prob}(S)$ the set of finitely supported probability measures on $S$. This is a semi-group under the operation of convolution. We say that $\Phi\colon\mathrm{Prob}(S)\to\mathrm{Prob}(S)$ is support-preserving if $\mu$ and $\Phi[\mu]$ are mutually absolutely continuous for all $\mu\in\mathrm{Prob}(S)$; note that since these are finitely supported measures, the two notions indeed coincide. We say that $\Phi$ is an endomorphism if it commutes with convolution, i.e., if $\Phi[\mu_{1}*\mu_{2}]=\Phi[\mu_{1}]*\Phi[\mu_{2}]$.
@@
is easily verified to be a support preserving endomorphism. Our main result is that these are the unique ones, when $S$ is $\mathbb{Z}_{\geq 0}$, $\mathbb{Z}$ or $\mathbb{Q}$.
@@
A corollary of Theorem 2 is that if $p$ is a polynomial with at least two terms—e.g., $p(x)=(x+1)/2$—then $\Phi$ is completely determined by $\Phi[p]$: if $\Phi^{\prime}[p]=\Phi[p]$ for two support-preserving multiplicative maps $\Phi,\Phi^{\prime}$, then $\Phi=\Phi^{\prime}$.
@@
=$\mathcal{M}=\left\{p(x)=\sum_{k=1}^{d}p_{k}x^{k}\,\middle|\,p(x)>0\text{ for all }x>0,\ p(1)=1\right\}.$
@@
Recall that, in the physics interpretation, a measure $\mu$ represents the mass of states at each energy level. In quantum mechanical settings, there are often only finitely many states at each energy level. In this case, the measure $\mu$ will have rational probabilities, and the corresponding probability-generating function will be a polynomial with rational coefficients. Indeed, if there are $k_{i}$ states with energy $E_{i}$ and the total number of states is $k=\sum_{i}k_{i}$, then $\mu$ has an atom of weight $k_{i}/k$ at $E_{i}$. This motivates the pursuit of the same questions, but in a rational setting. Let
@@
be the set of polynomials with non-negative rational coefficients that sum to one. As in the real case, a map $\Phi\colon\mathcal{P}_{\mathbb{Q}}\to\mathcal{P}$ is said to be support-preserving and multiplicative if $\Phi[\mu]$ and $\mu$ are mutually absolutely continuous and if $\Phi[\mu_{1}*\mu_{2}]=\Phi[\mu_{1}]*\Phi[\mu_{2}]$. The next result shows that Theorem 2 still holds in this setting.
@@
is a support-preserving endomorphism. Of course, this is non-constructive, since the existence of non-continuous solutions of the Cauchy equation requires an application of some axiom of choice (e.g., any axiom that is strong enough to guarantee a Hamel basis of $\mathbb{R}$ over $\mathbb{Q}$). A natural conjecture is that any support preserving endomorphism on $\mathrm{Prob}(\mathbb{R})$ that is not continuous (equivalently, not of the form $\Phi_{\beta}$) is not measurable.
@@
Support preserving endomorphisms of $\mathrm{Prob}(\mathbb{Z}^{d})$ and beyond
@@
More generally, given a semigroup $G$, does there always exist a homomorphism $\pi\colon G\to\mathbb{R}$ such that every support-preserving $\Phi\colon\mathrm{Prob}(G)\to\mathrm{Prob}(G)$ is of the form
@@
For $S\subset\mathbb{R}$, say that $\Psi\colon\mathrm{Prob}(S)\to\mathrm{Prob}(S)$ is weakly support-preserving if $\Psi[\mu]$ is absolutely continuous with respect to $\mu$, i.e., the support of $\Psi[\mu]$ is a subset of the support of $\mu$. Clearly, every support-preserving $\Phi$ is also weakly support-preserving, and so the class of support-preserving endomorphisms of, say, $\mathrm{Prob}(\mathbb{Z}_{\geq 0})$ is contained in the weakly support-preserving ones. We conjecture that the set of all weakly support-preserving endomorphisms is exhausted by $\Phi_{\beta}$ and the two limiting cases $\Phi_{+\infty}=\lim_{\beta\to+\infty}\Phi_{\beta}$ and $\Phi_{-\infty}=\lim_{\beta\to-\infty}\Phi_{\beta}$, which correspond to putting a point mass on the maximal or minimal point of the support, respectively.
@@
The algebra of the semi-group of probability measures on the reals under convolutions is well studied, including its homomorphisms to $\mathbb{R}$, $\mathbb{Z}$ and $\mathbb{C}$; see “Algebraic Probability Theory,” a book by Ruzsa and Székely [ruzsa1988algebraic], as well as more recent work [mu2021monotone, fritz2019monotone]. Homomorphisms into general groups were considered by Mattner [mattner2004cumulants].
@@ 2. Preliminaries
$\Psi_{\gamma}\colon\mathcal{M}$ $\to\mathcal{M}$
@@
Note that we here again drop the normalization constant and write $(x+1)^{n}$ rather than $2^{-n}(x+1)^{n}$. This result is a consequence of Pólya’s Positivstellensatz [polya1928uber], a more general statement about multivariate polynomials; see also Theorem 56 in [hardy1952inequalities]. In the univariate setting, a version of this result was obtained by Bàlinth; see a footnote on the first page of [polya1928uber].
@@ 3.1. Extending the domain of $\Phi$
Hence $\Phi[x+a]\neq\Phi[x+b]$. We have thus proved the following claim: [^5] Again, we omit normalization constants as discussed in §2. With these constants the claim would be that $a\neq b$ implies $\Phi[(x+a)/(1+a)]\neq\Phi[(x+b)/1+b)]$.
@@
Let $\Phi\colon\mathcal{P}\to\mathcal{P}$ be support preserving and multiplicative. Extend it to $\Phi\colon\mathcal{P}\to\mathcal{F}(\mathbb{Z}_{\geq 0})$ using (3.1). Choose any $a,b>0$, $a\neq b$, and fix $q\in\mathcal{M}$. By Lemma 2.2, there is $n$ large enough so that both $q(x)(x+a)^{n}$ and $q(x)(x+b)^{n}$ are in $\mathcal{P}$. We thus have that
@@
Both sides are polynomials and have the same roots. Of these roots, at least $n$ are equal to $-c$, since the right-hand side includes the term $(x+c)^{n}$. The left-hand side thus also has at least $n$ roots which are equal to $-c$. Since $c\neq d$, $\Phi[q(x)\cdot(x+a)^{n}]$ has at least $n$ roots equal to $-c$. It follows that
@@ 3.2. Linear polynomials and some quadratic polynomials
$0\displaystyle 2\varphi(a)\varphi(b)\varphi(2b)-\varphi(a)\varphi(2b)^{2}=$ 0
@@
Consider an extension of a support-preserving and multiplicative $\Phi\colon\mathcal{P}\to\mathcal{P}$ to $\mathcal{M}$ and define $\gamma$ by $\Phi[x+1]=x+\gamma$. Then $\Phi[x^{2}-ax+ba]=x-\gamma ax+\gamma^{2}ab$ for all $0<a<b$.
@@ 3.3. Quadratic polynomials
the coefficients of $x$ and $x^{2}$ must vanish. We obtain $01-a^{\prime}=$ 0 and $b^{\prime}-a^{\prime}=0$. Thus $a^{\prime}=b^{\prime}=1$ and so $\Phi[q^{a}]=q^{a}$ for $a=1$.
@@
Let us show that $\gamma=1$. Observe that for $a=-1$
@@
Let $A$ be a subset of $(-2,2)$. Suppose that $\Phi^{\pi}[q^{a}]=q^{a}$ for all $a\in A$. Then $\Phi[q^{a}]=q^{a}$ for all $a$ such that $a^{2}-2\in A$.
@@
If a support-preserving multiplicative $\Phi$ satisfies $\Phi[x+1]=x+1$, then $\Phi[q^{a}]=q^{a}$ for all $a$ in $(-2,2)$.
@@
Let $A^{*}$ be the set of all $a\in(-2,2)$ such that $\Phi^{\prime}[q^{a}]=q^{a}$ for all support-preserving multiplicative maps $\Phi^{\prime}$ such that $\Phi^{\prime}[x+1]=x+1$. By Claim 3.8, $[-1,1]\subseteq A^{*}$.
Let $f(x)=x^{2}-2$, and denote by $f^{(n)}$ the $n$-fold composition of $f$ with itself. We claim that for any $x\in(1,2)$ there is a number $n$ such that $f^{(n)}(x)\in[-1,1]$. Since the image of $(1,2)$ under $f$ is $(-1,2)$, it is enough to show that there is no $x_{0}\in(1,2)$ such that $x_{n}=f^{(n)}(x_{0})$ stays in $(1,2)$ for all $n$. Towards a contradiction, suppose that such $x_{0}$ exists. For $x\in(1,2)$, we have that $f(x)<3x-4$, since $f(1)=-1$, $f(2)=2$ and $f$ is strictly convex. In particular, $f(x)-x<2x-4<0$ for all $x\in[1,2)$, so that $f(x)<x$. Thus the sequence $x_{n}$ is decreasing. Denote $x_{\infty}=\lim_{n}x_{n}\in[1,2)$. By continuity of $f$, we get $f(x_{\infty})=x_{\infty}$. But $f(x)<x$ for all $x\in[1,2)$. This contradiction implies that for any $x\in(1,2)$, there is $n$ such that $f^{(n)}(x)\in(-1,1)$. The same argument applies to $x\in(-2,-1)$, since $f(-x)=f(x)$.
By Claim 3.9, if $f(a)\in A^{*}$ then $a\in A^{*}$. It follows that if $f^{(n)}(a)\in A^{*}$, then $a\in A^{*}$. Since $[-1,1]\subseteq A^{*}$, we conclude that $(-2,2)\subseteq A^{*}$. Thus $A^{*}=(-2,2)$.
@@ 3.4. Proofs of Theorems 1 and 2
We first claim that to prove the theorem, it suffices to show that a support-preserving and multiplicative $\Phi$ such $\Phi[x+1]=x+1$ is the identity map. To see that this statement implies the theorem, let $\Phi$ be support-preserving and multiplicative. Define $\gamma>0$ by $\Phi[x+1]=x+\gamma$. Recall from (2) that $\Psi_{\gamma}\colon\mathcal{P}\to\mathcal{P}$ is the map that takes $p(x)$ to $p(x/\gamma)$. Hence $\Phi^{\prime}=\Psi_{\gamma}^{-1}\circ\Phi$ is support-preserving and multiplicative, and furthermore satisfies $\Phi^{\prime}[x+1]=x+1$. Hence, if we show that $\Phi^{\prime}$ is the identity map, then we have shown that $\Phi=\Psi_{\gamma}$.

### v3

# On the origin of the Boltzmann Distribution

@@ Abstract.
The family of Boltzmann distributions is used in statistical mechanics to describe the distribution of states in systems with a given temperature. We give a novel characterization of this family as the unique one satisfying independence for uncoupled systems. The theorem boils down to a statement about endomorphisms of the convolution semi-group of finitely supported probability measures on the natural numbers, or, alternatively, about endomorphisms of the multiplicative semi-group of polynomials with non-negative coefficients.
@@ 1. Introduction
In statistical mechanics, a system can be found in various possible states. A state refers to a complete microscopic description of the system which encodes all physically relevant variables, and in particular determines the system’s energy. [^1] Statistical mechanics serves only as a motivation for the mathematical question we study. For a brief introduction to the relevant physics concepts, see Section 2.1 of [mezard2009information]. An uninterested reader may proceed directly to the “Definitions and Results” section. In an equilibrium with a given temperature, the distribution of states is given by the Boltzmann distribution: The probability that the system has energy $E$ is proportional to the number of states with energy $E$ times $\mathrm{e}^{-\beta E}$, where $1/\beta$ is proportional to the temperature of the environment. Formally, if $\mu$ is a probability measure on the reals describing the fraction of states with given energy—i.e., for measurable $A\subseteq\mathbb{R}$, the fraction of states with energy in $A$ is $\mu(A)$—then the distribution of energy is given by the measure $\Phi_{\beta}[\mu]$ where
@@
In terms of the physics, the first property means that the system can only be found at an energy level that corresponds to an existing state, and conversely, any existing state can be attained. The second property involves products of independent systems. Suppose that $\mu_{1}$ and $\mu_{2}$ describe the distribution of states in two systems. Form a new system whose set of states is the product of the two sets of states, and whose energy in each state is the sum of the two corresponding energies; this corresponds to no interaction between the two subsystems. Then the convolution $\mu_{1}*\mu_{2}$ describes the new system, and (1.1) follows from the assumption that the joint distribution of states is the product measure of the distributions in the two subsystems, or, differently put, the assumption that there is no correlation between systems that do not interact.
@@
Given a subset $S\subseteq\mathbb{R}$ closed with respect to addition, denote by $\mathrm{Prob}(S)$ the set of finitely supported probability measures on $S$. This is a semi-group under the operation of convolution. We say that $\Phi\colon\mathrm{Prob}(S)\to\mathrm{Prob}(S)$ is support-preserving if $\mu$ and $\Phi[\mu]$ have the same support, or, equivalently, are mutually absolutely continuous, for all $\mu\in\mathrm{Prob}(S)$; note that since these are finitely supported measures, the two notions indeed coincide. We say that $\Phi$ is an endomorphism if it commutes with convolution, i.e., if $\Phi[\mu_{1}*\mu_{2}]=\Phi[\mu_{1}]*\Phi[\mu_{2}]$.
@@
is easily verified to be a support-preserving endomorphism. Our main result is that these are the unique ones, when $S$ is $\mathbb{Z}_{\geq 0}$, $\mathbb{Z}$ or $\mathbb{Q}$.
@@
A corollary of Theorem 2 is that if $p$ is a polynomial with at least two terms—e.g., p(x)=(x+1)/2p(x)=(x+1)/2—then $\Phi$ is completely determined by $\Phi[p]$: if $\Phi^{\prime}[p]=\Phi[p]$ for two support-preserving multiplicative maps $\Phi,\Phi^{\prime}$, then $\Phi=\Phi^{\prime}$.
@@
$\mathcal{M}=\left\{p(x)=\sum_{k=0}^{d}p_{k}x^{k}\,\middle|\,p(x)>0\text{ for all }x>0,\ p(1)=1\right\}.$
@@
Recall that, in the physics interpretation, a measure $\mu$ represents the fraction of states at each energy level. In quantum mechanical settings, there are often only finitely many states at each energy level. In this case, the measure $\mu$ will have rational probabilities, and the corresponding probability-generating function will be a polynomial with rational coefficients. Indeed, if there are $k_{i}$ states with energy $E_{i}$ and the total number of states is $k=\sum_{i}k_{i}$, then $\mu$ has an atom of weight $k_{i}/k$ at $E_{i}$. This motivates the pursuit of the same questions, but in a rational setting. Let
@@
be the set of polynomials with non-negative rational coefficients that sum to one. The next result shows that Theorem 2 still holds in this setting.
@@
is a support-preserving endomorphism. Of course, this is non-constructive, since the existence of non-continuous solutions of the Cauchy equation requires an application of some axiom of choice (e.g., any axiom that is strong enough to guarantee a Hamel basis of $\mathbb{R}$ over $\mathbb{Q}$). A natural conjecture is that any support-preserving endomorphism on $\mathrm{Prob}(\mathbb{R})$ that is not continuous (equivalently, not of the form $\Phi_{\beta}$) is not measurable.
@@
Support-preserving endomorphisms of $\mathrm{Prob}(\mathbb{Z}^{d})$ and beyond
@@
More generally, given a semi-group $G$, does there always exist a homomorphism $\pi\colon G\to\mathbb{R}$ such that every support-preserving $\Phi\colon\mathrm{Prob}(G)\to\mathrm{Prob}(G)$ is of the form
@@
For $S\subset\mathbb{R}$, say that $\Psi\colon\mathrm{Prob}(S)\to\mathrm{Prob}(S)$ is weakly support-preserving if $\Psi[\mu]$ is absolutely continuous with respect to $\mu$, i.e., the support of $\Psi[\mu]$ is a subset of the support of $\mu$. Clearly, every support-preserving $\Phi$ is also weakly support-preserving, and so the class of support-preserving endomorphisms of, say, $\mathrm{Prob}(\mathbb{Z}_{\geq 0})$ is contained in the weakly support-preserving ones. We conjecture that the set of all weakly support-preserving endomorphisms is exhausted by $\Phi_{\beta}$ and the two limiting cases $\Phi_{+\infty}=\lim_{\beta\to+\infty}\Phi_{\beta}$ and $\Phi_{-\infty}=\lim_{\beta\to-\infty}\Phi_{\beta}$, which correspond to putting a point mass on the minimal or maximal point of the support, respectively.
@@
The algebra of the semi-group of probability measures on the reals under convolutions is well studied, including its homomorphisms to $\mathbb{R}$, $\mathbb{Z}$ and $\mathbb{C}$; see “Algebraic Probability Theory,” a book by Ruzsa and Székely [ruzsa1988algebraic], as well as more recent work [mu2024monotone, fritz2019monotone]. Homomorphisms into general groups were considered by Mattner [mattner2004cumulants].
@@ 2. Preliminaries
$\Psi_{\gamma}\colon\mathcal{M} \displaystyle\to\mathcal{M}$
@@
Note that we here again drop the normalization constant and write $(x+1)^{n}$ rather than $2^{-n}(x+1)^{n}$. This result is a consequence of Pólya’s Positivstellensatz [polya1928uber], a more general statement about multivariate polynomials; see also Theorem 56 in [hardy1952inequalities]. In the univariate setting, a version of this result was obtained by Bálinth; see a footnote on the first page of [polya1928uber].
@@ 3.1. Extending the domain of $\Phi$
Hence $\Phi[x+a]\neq\Phi[x+b]$. We have thus proved the following claim: [^5] Again, we omit normalization constants as discussed in §2. With these constants the claim would be that $a\neq b$ implies $\Phi[(x+a)/(1+a)]\neq\Phi[(x+b)/(1+b)]$.
@@
Let $\Phi\colon\mathcal{P}\to\mathcal{P}$ be support-preserving and multiplicative. Extend it to $\Phi\colon\mathcal{P}\to\mathcal{F}(\mathbb{Z}_{\geq 0})$ using (3.1). Choose any $a,b>0$, $a\neq b$, and fix $q\in\mathcal{M}$. By Lemma 2.2, there is $n$ large enough so that both $q(x)(x+a)^{n}$ and $q(x)(x+b)^{n}$ are in $\mathcal{P}$. We thus have that
@@
Both sides are polynomials and have the same roots. Of these roots, at least $n$ are equal to c-c, since the right-hand side includes the term $(x+c)^{n}$. The left-hand side thus also has at least $n$ roots which are equal to c-c. Since $c\neq d$, $\Phi[q(x)\cdot(x+a)^{n}]$ has at least $n$ roots equal to c-c. It follows that
@@ 3.2. Linear polynomials and some quadratic polynomials
$2\varphi(a)\varphi(b)\varphi(2b)-\varphi(a)\varphi(2b)^{2}=0$
@@
Consider an extension of a support-preserving and multiplicative $\Phi\colon\mathcal{P}\to\mathcal{P}$ to $\mathcal{M}$ and define $\gamma$ by $\Phi[x+1]=x+\gamma$. Then $\Phi[x^{2}-ax+ba]=x^{2}-\gamma ax+\gamma^{2}ab$ for all $0<a<b$.
@@ 3.3. Quadratic polynomials
the coefficients of $x$ and $x^{2}$ must vanish. We obtain $1-a^{\prime}=0$ and $b^{\prime}-a^{\prime}=0$. Thus $a^{\prime}=b^{\prime}=1$ and so $\Phi[q^{a}]=q^{a}$ for $a=1$.
@@
Let us show that $\gamma=1$. Observe that for a=1a=-1
@@
Let $A$ be a subset of (2,2)(-2,2). Suppose that $\Phi^{\pi}[q^{a}]=q^{a}$ for all $a\in A$. Then $\Phi[q^{a}]=q^{a}$ for all $a$ such that $a^{2}-2\in A$.
@@
If a support-preserving multiplicative $\Phi$ satisfies $\Phi[x+1]=x+1$, then $\Phi[q^{a}]=q^{a}$ for all $a$ in (2,2)(-2,2).
@@
Let $A^{*}$ be the set of all $a\in(-2,2)$ such that $\Phi^{\prime}[q^{a}]=q^{a}$ for all support-preserving multiplicative maps $\Phi^{\prime}$ such that $\Phi^{\prime}[x+1]=x+1$. By Claim 3.8, $1,1]\subseteq A^{*}$.
Let $f(x)=x^{2}-2$, and denote by $f^{(n)}$ the $n$-fold composition of $f$ with itself. We claim that for any $x\in(1,2)$ there is a number $n$ such that $f^{(n)}(x)\in[-1,1]$. Since the image of $(1,2)$ under $f$ is (1,2)(-1,2), it is enough to show that there is no $x_{0}\in(1,2)$ such that $x_{n}=f^{(n)}(x_{0})$ stays in $(1,2)$ for all $n$. Towards a contradiction, suppose that such $x_{0}$ exists. For $x\in(1,2)$, we have that f(x)<3x4f(x)<3x-4, since f(1)=1f(1)=-1, $f(2)=2$ and $f$ is strictly convex. In particular, f(x)x<2x4<0f(x)-x<2x-4<0 for all $x\in[1,2)$, so that $f(x)<x$. Thus the sequence $x_{n}$ is decreasing. Denote $x_{\infty}=\lim_{n}x_{n}\in[1,2)$. By continuity of $f$, we get $f(x_{\infty})=x_{\infty}$. But $f(x)<x$ for all $x\in[1,2)$. This contradiction implies that for any $x\in(1,2)$, there is $n$ such that $f^{(n)}(x)\in(-1,1)$. The same argument applies to $x\in(-2,-1)$, since f(x)=f(x)f(-x)=f(x).
By Claim 3.9, if $f(a)\in A^{*}$ then $a\in A^{*}$. It follows that if $f^{(n)}(a)\in A^{*}$, then $a\in A^{*}$. Since $1,1]\subseteq A^{*}$, we conclude that $(-2,2)\subseteq A^{*}$. Thus $A^{*}=(-2,2)$.
@@ 3.4. Proofs of Theorems 1 and 2
We first claim that to prove the theorem, it suffices to show that a support-preserving and multiplicative $\Phi$ such $\Phi[x+1]=x+1$ is the identity map. To see that this statement implies the theorem, let $\Phi$ be support-preserving and multiplicative. Define $\gamma>0$ by $\Phi[x+1]=x+\gamma$. Recall from (2) that $\Psi_{\gamma}\colon\mathcal{P}\to\mathcal{P}$ is the map that takes $p(x)$ to $p(x/\gamma)/p(1/\gamma)$. Hence $\Phi^{\prime}=\Psi_{\gamma}^{-1}\circ\Phi$ is support-preserving and multiplicative, and furthermore satisfies $\Phi^{\prime}[x+1]=x+1$. Hence, if we show that $\Phi^{\prime}$ is the identity map, then we have shown that $\Phi=\Psi_{\gamma}$.